# Defining Boundaries: The Impact of Domain Specification on Cross-Language and Cross-Domain Transfer in Machine Translation

### v1

@@ Abstract
Recent advancements in neural machine translation (NMT) have revolutionized the field, yet the dependency on extensive parallel corpora limits progress for low-resource languages. Cross-lingual transfer learning offers a promising solution by utilizing data from high-resource languages but often struggles with in-domain NMT. In this paper, we investigate three pivotal aspects: enhancing the domain-specific quality of NMT by fine-tuning domain-relevant data from different language pairs, identifying which domains are transferable in zero-shot scenarios, and assessing the impact of language-specific versus domain-specific factors on adaptation effectiveness. Using English as the source language and Spanish for fine-tuning, we evaluate multiple target languages including Portuguese, Italian, French, Czech, Polish, and Greek. Our findings reveal significant improvements in domain-specific translation quality, especially in specialized fields such as medical, legal, and IT, underscoring the importance of well-defined domain data and transparency of the experiment setup in in-domain transfer learning.
@@ 2.3 Zero-Shot Cross-Lingual Domain Adaptation
Building on transfer learning across languages and domain adaptation, zero-shot cross-lingual domain adaptation tackles adapting multilingual NMT to specialized domains for languages with limited parallel in-domain data. The approach leverages a multilingual pre-trained NMT model fine-tuned on domain-specific data from a high-resource language pair, enabling it to capture domain knowledge. This adapted model can then translate target languages within the same domain, transferring domain knowledge in a zero-shot manner. The effectiveness of this approach depends on several influencing factors, including the linguistic proximity between the pivot and target languages, the nature and complexity of the domain itself, as well as the composition of the initial general-purpose pre-training data. Moreover, as the model transitions from general pre-training to specialized fine-tuning, there is also a risk of catastrophic forgetting where previously learned general knowledge is overwritten (Saunders, 2022). Approaches like embedding freezing (Grosso et al., 2022) have proved effective in mitigating this issue. While Grosso et al. (2022) demonstrated the feasibility of zero-shot cross-lingual domain adaptation for the medical domain, comprehensive analysis across diverse languages and domains is still needed to fully understand these influencing factors and their relative impacts.

### v2

@@ Abstract
Recent advancements in neural machine translation (NMT) have revolutionized the field, yet the dependency on extensive parallel corpora limits progress for low-resource languages and domains. Cross-lingual transfer learning offers a promising solution by utilizing data from high-resource languages but often struggles with in-domain NMT. This paper investigates zero-shot cross-lingual domain adaptation for NMT, focusing on the impact of domain specification and linguistic factors on transfer effectiveness. Using English as the source language and Spanish for fine-tuning, we evaluate multiple target languages, including Portuguese, Italian, French, Czech, Polish, and Greek. We demonstrate that both language-specific and domain-specific factors influence transfer effectiveness, with domain characteristics playing a crucial role in determining cross-domain transfer potential. We also explore the feasibility of zero-shot cross-lingual cross-domain transfer, providing insights into which domains are more responsive to transfer and why. Our results show the importance of well-defined domain boundaries and transparency in experimental setups for in-domain transfer learning.
@@ 2.3 Zero-Shot Cross-Lingual Domain Adaptation
Building on transfer learning across languages and domain adaptation, zero-shot cross-lingual domain adaptation tackles adapting multilingual NMT to specialized domains for languages with limited parallel in-domain data. The approach leverages a multilingual pre-trained NMT model fine-tuned on domain-specific data from a high-resource language pair, enabling it to capture domain knowledge. This adapted model can then translate target languages within the same domain, transferring domain knowledge in a zero-shot manner. The effectiveness of this approach depends on several influencing factors, including the linguistic proximity between the pivot and target languages, the nature and complexity of the domain itself, as well as the composition of the initial general-purpose pre-training data. Moreover, as the model transitions from general pre-training to specialized fine-tuning, there is also a risk of catastrophic forgetting where previously learned general knowledge is overwritten (Saunders, 2022). Approaches like embedding freezing (Grosso et al., 2022) have proved effective in mitigating this issue. While Grosso et al. (2022) demonstrated the feasibility of zero-shot cross-lingual domain adaptation for the medical domain, comprehensive analysis across diverse languages and domains is still needed to understand these influencing factors and their relative impacts fully.